OpenStudy (anonymous):

What is the equation of a line that passes through the point (– 3,–11) and is parallel to the line whose equation is 2x – y = 4? (please use detail)

OpenStudy (anonymous):

This equation is in slope-intercept form, y=mx+b. m is the slope of the line, and b is the y-intercept. Start out with that. On the y-axis, make a point at -4. Since the slope is positive 2, and slope is rise over run, you can move up two and one to the right, and make another point. Do that until you can draw a straight line through (you really only need two points), and that is the first equation. To graph the inverse, you need to find the equation first. To do this, you have to switch x and y. So: y=2x-4 (switch x and y) x=2y-4 (Now solve for y to get it back into y=mx+b form) You get: 2y=x+4 Then: y=x/2 +2 To graph that, you do the same thing. The y-intercept is now +2, so make a point there on the y-axis. The slope is now 1/2, so from your first point, move up one and two to the right. Draw a line through the points, and that's the second graph. For the points of intersection, if you have a good graph you can probably just eyeball it. To do it algebraically though, you need to set the equations equal to one another. So, if you have y=2x-4 and y=x/2+2, do: 2x-4 = x/2+2 (Solve for x) 2x = x/2+6 4x = x+12 3x=12 x=4 (Now plug this into one of your first equations to get a y-value) y=2(4)-4 y=4 So your POI is (4,4)

OpenStudy (anonymous):

that's not one of my options I don't think! I'm doing some multiple choice! and it's

OpenStudy (anonymous):

[1] y = 2x + 5

OpenStudy (anonymous):

2] y = 2x - 5

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

Just those two?

OpenStudy (anonymous):

noo there's another

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

3. y=1/2 x + 25/2 4. y=-1/2x-25/2

OpenStudy (anonymous):

this is for geometry btw haha

OpenStudy (anonymous):

Given the equation of the line, 2x – y = 4 Rearrange to the general form of y=mx+c 2x-y = 4 y= 2x-4 Gradient of the line is the coefficient of x, which is 2 in this case. Given the equation of the line you want to find is y=mx+c Substitute m=2 and the point (-3,-11) to y=mx+c y=mx+c -11 = 2(-3) +c -11= -6 +c c= -5 Eqn. of line = y= 2x-5

OpenStudy (anonymous):

wow thanks!:)